OpenStudy (anonymous):

-x + 3y = -9 by graphing the x and y intercepts

OpenStudy (johnweldon1993):

So to find the y-intercept you need to set x = to 0 in your equation...and solve for y And to find the x-intercept...set y = to 0 and solve for x So y-intercept: -x + 3y = -9 -0 + 3y = -9 3y = -9 y = ? x-intercept: -x + 3y = -9 -x + 3(0) = -9 -x + 0 = -9 -x = -9 x = ?